A two-digit number is the same as the number with its digits reversed increased by three times the original number's unit's digit. If five times the sum of the digits is 70, find the original number. (find a system of two equations then use the system to find the original number).
Answer by checkley77(12569)

10x+y=10y+x+3y
10x-x+y-10y-3y=0
9x-12y=0 multiply by -5
---------------
5(x+y)=70
5x+5y=70 multiply by 9 & add
---------------
-45x+60y=0
45x+45y=630
-----------
105y=630
y=630/105
y=6 answer for the tens digit in the original number.
5(x+6)=70
5x+30=70
5x=70-30
5x=40
x=40/5
x=8
Proof:
10*8+6=10*6+8+3*6
80+6=60+8+18
86=86